OpenStudy (countrygirl16):

10th grade geometry hell please!!?

OpenStudy (blurbendy):

do you have a question?

OpenStudy (countrygirl16):

here one if u could plz jus explain how i get the answer that would be great!!

OpenStudy (countrygirl16):

i typed help lol somehow it put h*ll o.O

OpenStudy (blurbendy):

wait i was gonna help!

OpenStudy (countrygirl16):

sorry lol thought u left :P so whats the first step to findin the answer?

OpenStudy (blurbendy):

ok so first you need to find the difference (use absolute value) between X1 and X2, so you get |2 - 18| = 16 Then the same for y |10 + 18| = 28 Now square both and add 16^2 + 28^2 = 1040 Then take the square root of 1040 That will give you the distance.

OpenStudy (countrygirl16):

So would the answer be 1081600? how do i round this lol its a huge number

OpenStudy (blurbendy):

\[\sqrt{1040} = \] you should get like 30 something

OpenStudy (countrygirl16):

oh! my bad fergot you have to type square root first on the calculator i got 32. a bunch of numbers so rounded would be 30? :D

OpenStudy (blurbendy):

nearest hundredth would mean you would round it to 32.25

OpenStudy (countrygirl16):

oh ok, so then that would be my answer right?

OpenStudy (blurbendy):

yup!

OpenStudy (countrygirl16):

:D thank you soooo much!

OpenStudy (blurbendy):

np! if you have any other questions just ask

OpenStudy (countrygirl16):

ok and yea i do hang on a sec

OpenStudy (blurbendy):

k

OpenStudy (countrygirl16):

alrighty how do i figure this one out?

OpenStudy (blurbendy):

hang on

OpenStudy (countrygirl16):

ok

OpenStudy (blurbendy):

okay, it's 11.7 sry it took so long.

OpenStudy (countrygirl16):

thanks :D can u tell me how u got it? that way i can learn to do this as well haha got a seg test comin up and nah thats ok

OpenStudy (blurbendy):

ya 1 sec

OpenStudy (blurbendy):

14 cos 60 + sqrt ( 13^2 - (14 cos 30)^2) = 11.7 Had to use the formula for the ASA theorem.

OpenStudy (countrygirl16):

ok makes since, thats all for tonight, thanks again :)

OpenStudy (countrygirl16):

sense*

OpenStudy (blurbendy):

no problem! just message me if you need help again.